OpenStudy (anonymous):

If 5^x = 7, what does 5^(-2x) equal

OpenStudy (anonymous):

i think 1/32

OpenStudy (anonymous):

see 5^x = 7 ( given ) 5^(-2x) can also be written as 5^(x)^(-2) = 7(-2) now as we know a^-b = 1/a^(b) so 7^-2 = 1/7^2 = 1/49

OpenStudy (anonymous):

give me medal